# Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data

### v1

@@ 6.2.1. Object Detection Experiments
((c)) BDD
Figure 5. Distribution of Average Precision for different queries for CLIP and Needle, split into hard and easy subsets.
In these experiments, we demonstrate the effectiveness of Needle in basic object detection tasks. We utilize state-of-the-art benchmark datasets, LVIS (Gupta et al., 2019), Caltech256 (Griffin et al., 2022), and BDD (Yu et al., 2020), for this evaluation. For each query, the number of retrieved images is set to 60 ($k=60$), and we aim to identify 10 relevant tuples within the results ($R=10$). Two local Stable Diffusion models are employed as Foundation Models: runwayml-v1.5 and sdxl-turbo-v1.0 with guide image resolution of $768\times 768$. For each model, two guide images are generated, resulting in a total of four guide images per query ($m=4$). Table 1 presents a comparative analysis of performance between CLIP and Needle across the different datasets. Notably, Needle consistently demonstrates superior performance across all categories, in particular, for the hard subsets, Needle achieves around 200% improvement in AP compared to CLIP on average, highlighting its effectiveness in addressing more challenging object detection scenarios. Figure 5 shows the AP distribution of CLIP and Needle for different datasets. Needle demonstrates significant improvements in AP, even for queries where CLIP initially performs very poorly, with AP close to zero.
It is important to note that all generated guide images are produced using local Stable Diffusion models, which generate lower quality images than commercial alternatives. This result highlights that Needle imposes much looser constraints on the quality of guide tuples than humans do. Even when the guide images are distorted or lack realism, Needle can robustly retrieve relevant images using these guide tuples.
6.2.2. Targeted Queries
Figure 6. Illustration of different specific queries, the generated guide images and the comparison of CLIP and Needle results
Table 2. Performance (mAP, mHR) Comparison Between CLIP and Needle on Targeted Queries Benchmark Datasets
Dataset (#queries) mAP mHR
CLIP Needle CLIP Needle
Nocaps (100) 0.11 0.37 0.55 0.73
SentiCap (100) 0.24 0.34 0.71 0.77
Average 0.18 0.36 0.63 0.75
Needle excels at handling targeted queries and identifying rare tuples. We demonstrate this by running two image captioning benchmarks, Nocaps (Agrawal et al., 2019), and SentiCap (Mathews et al., 2016), using both CLIP and Needle. For each experiment, the systems are provided with a query that describes a specific image in the dataset, with the search engines returning 10 images. For this task, we enhance Needle with two more powerful and precise Foundation Models: DALL·E 3  [^4] https://openai.com/index/dall-e-3/ and Flux-Schnell  [^5] https://replicate.com/black-forest-labs/flux-schnell. These models are selected due to the need for guide images that accurately represent the detailed input queries, as these queries include significant complexity and detail.
In addition to reporting mean Average Precision (mAP), we also evaluate the mean Hit Rate (mHR). Hit Rate is particularly useful in cases where the number of retrieved images is small, and the focus is on finding the correct image rather than its specific ranking.
As shown in Table 2, Needle significantly outperforms CLIP in locating the target image. Needle successfully retrieves the correct image in over 75% of queries within the top 10 results, marking around 20% improvement over CLIP’s performance in mHR and 100% improvement in mAP. Figure 6 presents four different queries from the Nocaps (Agrawal et al., 2019) dataset, along with the generated guide images for Needle and the top-1 results from both CLIP and Needle. As illustrated, CLIP frequently overlooks key aspects of the input prompt and focuses on a single keyword rather than capturing the full intent of the query. In contrast, Needle effectively captures all inherent complexity of query within its guide tuples and identifies the target image using only two guide images.
6.3. Parameter Tuning
((a)) Runwayml (Caltech256)
((b)) sdxl-turbo (Caltech256)
((c)) Runwayml (LVIS)
((d)) sdxl-turbo (LVIS)
Figure 7. Heatmaps of Needle mAP for two different datasets (Caltech256 and LVIS) using two local Foundation Models (Runwayml and sdxl-turbo) while altering number of Generated Images $m$ and Generated Image Size
In this section, we investigate the effect of varying the number of generated guide images ($m=\{1,2,4\}$), the image resolution ($\{512\times 512,768\times 768,1024\times 1024\}$), and the performance of different foundation models (runwayml, sdxl-turbo) on Needle’s effectiveness. The experiments are conducted on both the Caltech256 and LVIS datasets. Figure 7 illustrates the mAP values for different combinations of these parameters. As a general trend, increasing the number of images and the resolution improves mAP performance. However, there are deeper insights to consider: first, sdxl-turbo exhibits more consistent and slightly better performance across different image sizes compared to runwayml; second, while increasing the number of generated images $m$ improves results, the improvement diminishes and stabilizes after a certain point. A similar trend is observed for image sizes. These results suggest that Needle can maintain high performance even with a limited number of generated images and lower resolution inputs.
6.4. Feedback Mode
((a)) LVIS
((b)) BDD
Figure 8. Comparison of AP distribution for Needle on Feedback mode on and off
CLIPNeedleBothNeither002020202040404040606060608080808010010010010023.2323.2323.2323.2352.5252.5252.5252.5214.1514.1514.1514.1510.110.110.110.1Percentage (%)
Figure 9. Case study results
Table 3. Comparison of Needle’s performance (mAP) for different datasets with Feedback mode enabled (ON) and built-in Anomaly detection (OFF)
Dataset (#hard/#all) All Queries Hard Subset
OFF ON OFF ON
BDD (2/11) 0.82 0.84 0.22 0.25
LVIS (84/100) 0.28 0.30 0.19 0.21
Average 0.55 0.57 0.21 0.23
In this subsection, we investigate the impact of enabling the feedback procedure in Needle on mAP. A human evaluator was employed to verify the quality and alignment of the generated images with the input query prior to retrieving results. For evaluation, we used LVIS and BDD datasets, generating four tuples per query ($m=4$) with an image size of $768\times 768$, utilizing the runwayml foundation model. We compare the mAP of Needle’s built-in anomaly detection method (§ 5.2) to the performance when human feedback is incorporated.
Figure 8 presents the AP distribution for both baselines. As shown, there is a slight improvement when feedback mode is enabled, primarily raising queries with $AP\approx 0.4$ in LVIS and $AP\approx 0.8$ in BDD toward higher APs. This indicates that: (a) Needle’s built-in anomaly detection method performs effectively in practice, achieving results comparable to a human evaluator, and (b) human intervention is most beneficial for queries that already have a moderate AP, whereas queries for which foundation models fail to generate suitable images do not significantly improve with feedback alone.
6.5. Case Study
@@
Each human evaluator was asked to write five queries of their choice, with varying levels of complexity. For each query, we retrieved 20 relevant images from both CLIP and Needle, displaying the results side by side with randomized orientations for each query. Evaluators were then asked to select the more relevant results, choosing from four options: ”Left is better,” ”Right is better,” ”Both are good enough,” and ”Neither is good enough.”, then we recorded the engine behind that option as preferred engine.
@@
A total of 20 human evaluators participated in the study, contributing 100 queries. The queries ranged from simple ones like ”Celery stew” to more complex descriptions such as ”A person walking on the sidewalk of a river in Fall at Chicago.”, Figure 9 presents the results of this user study. If we consider only the queries where one engine is preferred over the other, Needle is chosen as the better engine in almost 70% of cases.
7. Related Work
Our paper relates to work in the following areas:
Multi-modal Data Retrieval: Multi-modal data retrieval, which integrates various data types (e.g., text, images, video, and audio), has attracted significant attention. Early methods focused on shared representations by mapping modalities into a common latent space using canonical correlation analysis (CCA) (Rasiwasia et al., 2010; Wang et al., 2015b) and kernel-based techniques (Jia et al., 2019). Recent approaches leverage deep learning models, including cross-modal encoders (Wang et al., 2015a; Liu et al., 2017) and joint embedding networks (Balaneshin-kordan and Kotov, 2018; Rubio et al., 2017). Pre-trained transformers like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) highlight the effectiveness of combining vision and language models for cross-modal tasks.
Advanced techniques, such as attention mechanisms (Yang et al., 2021; Tang et al., 2023) and contrastive learning (Pham et al., 2024), further enhance retrieval by capturing fine-grained relationships. Feedback mechanisms (Moll et al., 2023; Zhou and Huang, 2003) and user preferences (Zhang and Zhang, 2010; Chen et al., 2020) improve multi-modal search results. Research also extends to other modalities, such as text-video retrieval (Li et al., 2019; Markatopoulou et al., 2017; Torabi et al., 2016; Mithun et al., 2018; Miech et al., 2019; Luo et al., 2021) and text-audio retrieval (Xin et al., 2023; Chechik et al., 2008; Koepke et al., 2022). For a detailed survey of cross-modal retrieval methods and future directions, see (Li et al., 2023).
GenAI in Data Management: The intersection of Generative AI (GenAI) and data management has emerged as a promising area, driving innovations in how data is generated, managed, and utilized. GenAI models, such as GPT (Achiam et al., 2023) and Stable Diffusion (Rombach et al., 2022), have been employed for synthetic data generation to augment datasets (Tanaka and Aranha, 2019; Shahbazi et al., 2024), improve model training (Le et al., 2017; Tremblay et al., 2018), and address data scarcity or imbalance issues (Erfanian et al., 2024). In data cleaning and integration, these models assist with tasks like imputation (Villaizán-Vallelado et al., 2024), and anomaly detection by generating plausible data patterns (Ye et al., 2024). Recent works also explore the role of GenAI in automating query generation (Trummer, 2022, 2023; Li and Xie, 2024) and data summarization (Dar et al., 2024), enhancing the efficiency of database systems and analytics platforms.
Natural Language Query Processing: The evolution of natural language to structured query translation has advanced from rule-based systems to sophisticated large language models (LLMs). Early methods used template matching and predefined rules (Androutsopoulos et al., 1995; Li and Jagadish, 2014), which were inflexible and required frequent manual updates. Machine learning approaches introduced more adaptability by learning query patterns from annotated datasets (Ma et al., 2020; Zhang et al., 2019), though they struggled with complex language constructs.
Deep learning models, especially transformers like BERT (Devlin, 2018) and GPT (Brown, 2020), enabled semantic understanding and laid the foundation for text-to-SQL systems (Li et al., 2024). Modern LLMs, such as PaLM (Chowdhery et al., 2023) and GPT-4 (Achiam et al., 2023), generate context-aware queries across multiple domains with minimal supervision. Retrieval-augmented generation (RAG) further enhances query precision by leveraging external knowledge (Lewis et al., 2020). Ongoing efforts focus on fine-tuning LLMs to improve accuracy and efficiency in converting natural language to structured queries (Trummer, 2022, 2023; Li and Xie, 2024).
8. Discussions and Limitations
8.1. Reliance on Existing Models
Needle extensively relies on Foundation Models for guide tuple generation and on embedders for generating vector representations of images. Both these components have their own capabilities, limitations, and inherent biases. This reliance presents both strengths and weaknesses in Needle’s architecture. The strength lies in the ability to easily improve performance by upgrading the internal Foundation Models and embedders. However, the limitations of these models also become limitations for Needle. To mitigate this, we designed Needle to support multiple Foundation Models and embedders, enabling it to draw from a broader range of knowledge and reduce the impact of individual model limitations.
8.2. Efficiency
We assess the efficiency of Needle across two phases: Preprocessing and Inference.
Preprocessing: In the preprocessing phase, two highly parallelizable tasks are performed: tiling and embedding creation. Tiling is a fast operation, but it can linearly increase the time required for embedding creation. To improve time efficiency, one can increase the minimum tile width and height, thereby reducing the number of tiles. The time required for embedding creation can be significantly improved by utilizing multiple GPUs. Since the preprocessing phase is a one-time operation that can be easily parallelized and fine-tuned, it does not present a major concern for Needle’s overall performance.
Inference: Needle’s inference consists of three main sequential operations: (a) generating guide tuples, (b) k-NN lookup, and (c) ranking aggregation. For an input dataset of size $n$, retrieving $k(1\leq k\leq n)$ results with constants $m$ (number of guide images) and $l$ (number of embedders), (a) is a constant-time operation with respect to $n$ and $k$. Step (b)’s time complexity depends on the underlying vector store algorithm, and for HNSW data structure (Malkov and Yashunin, 2018), is $O(k\log(n))$. Step (c) has a time complexity of $O(k)$ using Fagin’s Algorithm (Fagin et al., 2001).
In our experiments, where the number of tuples (tiles) is at most 400,000 for any k, (b) k-NN lookup and (c) the rankings aggregation take less than one second using current approximate vector stores. The bottleneck of the entire process lies in (a), generating guide tuples. The good news is that this operation is independent of the dataset size $n$ and number of images to retrieve $k$, and only relies on the number of guide images $m$. Moreover, the generation process can run in parallel, significantly improving performance.
To summarize, for small values of $n$, Needle’s inference time may be slower than other methods due to the constant time required for guide image generation. However, as $n$ increases, the effect of step (a) diminishes, and Needle’s efficiency becomes comparable to other methods.
8.3. Extensions to Other Modalities
The core idea behind Needle is adaptable to various modalities beyond images, including audio, video, and others. However, due to the current limitations in publicly available foundation models for audio and video generation, our focus in this paper has been restricted to the image modality. As foundation models for these additional modalities continue to improve and become more accessible, we plan to extend Needle to support them. This will allow Needle to operate across multiple data types, unlocking further potential for handling complex, multi-modal queries in the near future.
9. Conclusion
Following this method, We implemented Needle for image retrieval. We proposed different solutions for practical challenges in our system’s development. Our comprehensive experiments demonstrated that our system, Needle, significantly outperforms the state-of-the-art techniques in image retrieval.
While our proposed technique is not limited to image datasets, its development for other modalities, such as audio and video, requires foundation models that are currently not publicly available. Upon the availability of such models, we aim to extend the scope of Needle to a wider range of modalities in our future work.

### v2

@@ E.4 Hyper-parameter Ablation Study
((c)) COCO
((d)) LVIS
Figure 7: Illustration of how varying number of guide images and number of embedders involved affects Needle MAP
In this section, we examine the impact of each hyper-parameter on Needle’s performance.
Generated Guide Images and Embedders Count: We investigate the impact of the number of images generated per engine ($m$) and the number of embedders used ($\ell$) on the performance of Needle. We vary $m$ over $\{1,2,3\}$ and $\ell$ over $\{1,2,4,6\}$, and conduct experiments with these configurations on the BDD, Caltech256, COCO, and LVIS datasets. While incrementing the value of $\ell$, we add embedders sorted by their weight in descending order (Higher performing embedders are being used first). Figure 7 shows the mean average precision (MAP) achieved for each combination. In general, increasing both $m$ and $\ell$ leads to improved MAP. Notably, even the lowest configuration ($m=1$, $\ell=1$) outperforms most of the baselines evaluated in Table 1 across different benchmarks.
From this ablation study, we conclude that consensus among embedders can lead to higher reliability and improved performance in Needle. Note that Needle supports an arbitrary number of embedders, allowing users to increase this number until performance saturates. This experiment also demonstrates that increasing the number of generated images enhances performance; specifically, the performance gap between $m=1$ and $m=2$ is larger than that between $m=2$ and $m=3$. This indicates that Needle can achieve good results with a small number of images and does not require extensive image generation to capture all relevant complexities.
Foundation Models: Needle supports image generation using various open-source and proprietary foundation models. These models differ in the quality of generated images, realism, and their alignment with the input prompt. In this section, we investigate two primary aspects. First, we assess the performance of each individual foundation model when used on its own. Second, we examine how combining multiple foundation models can improve performance by providing diverse representations of the input query.
Our experiments involve four foundation models: DALL-E3 (proprietary), Imagen3-fast (proprietary), Flux-Schnell (open-source), and RealViSXL (based on Stable-diffusionXL; open-source). We evaluate their performance on object detection tasks using the BDD, Caltech256, and COCO datasets, as well as on complex natural language queries using the Winoground and COLA benchmarks. The inclusion of tests on complex natural language queries is motivated by the higher demands for prompt alignment and image quality in such scenarios. We calculate MAP for Object detection and Accuracy for Complex natural language benchmarks as the performance metric. Figure 8 demonstrates how each foundation model performed in different benchmarks.
Notably, the two open-source models, Flux-Schnell and RealVisXL, outperformed proprietary models by a huge margin in Complex queries, indicating they have better prompt alignment and understanding compared to DALL-E3 and Imagen3-fast. DALL-E3 performed worse than others in most of the benchmarks, and RealVisXL demonstrated to be a reliable option for most cases.
Figure 9 illustrates that increasing the number of foundation models leads to an improvement in the overall performance of Needle. Notably, the addition of further foundation models yields more improvement in performance for complex natural language queries than for object detection benchmarks. This outcome is likely attributable to the increased likelihood of generating images that accurately capture the nuanced details of the input prompt, thereby facilitating the correct identification of the target image. In contrast, for object detection tasks—where the objects are generally simpler to generate—the primary benefit of including additional foundation models is an increase in recall due to the diverse representations of the target object across various formats and styles.
0.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111BDDCaltech256COCO0.690.690.690.690.910.910.910.910.960.960.960.960.730.730.730.730.920.920.920.920.950.950.950.950.610.610.610.610.910.910.910.910.930.930.930.930.680.680.680.680.940.940.940.940.960.960.960.96MAPObject Detection BenchmarksImagen3-fastFlux-schnellDALL-ERealVisXL0.40.40.40.40.450.450.450.450.50.50.50.50.550.550.550.550.60.60.60.60.650.650.650.65COLAWinoground0.60.60.60.60.430.430.430.430.620.620.620.620.510.510.510.510.610.610.610.610.440.440.440.440.610.610.610.610.530.530.530.53AccuracyComplex NLQ Benchmarks
Figure 8: Comparison of Performance of Different Image Generators on Object Detection and Complex Natural Language Query benchmarks.
11112222333344440.70.70.70.70.80.80.80.80.90.90.90.91111#Foundation ModelsMAPObject Detection BenchmarksBDDCOCOCaltech256
11112222333344440.50.50.50.50.550.550.550.550.60.60.60.60.650.650.650.650.70.70.70.7#Foundation ModelsAccuracyComplex NLQ BenchmarksWinogroundCOLA
Figure 9: The effect of number of Foundation Models on the Performance of Needle for different Benchmarks
SMALLMEDIUMLARGE0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8Image SizeMAPBDDDALL-EFlux-SchnellRealVisXLSMALLMEDIUMLARGE0.80.80.80.80.90.90.90.91111Image SizeCOCO
Figure 10: Illustration of variation of Needle MAP based on size of guide image for different foundation models
Image Quality: Foundation models usually support generating images in different resolutions. In this ablation study, we investigate the effect of image resolution (size) of guide images on Needle’s performance. We define three supported image sizes for each foundation model, SMALL, MEDIUM and LARGE. The exact amount of pixels is determined based on the foundation model design. We generate 3 images per query and utilize Flux-Schnell, RealVisXL and DALL-E for this study (ImagenV3-fast does not support different images sizes as input). Figure 10 illustrates the MAP performance of various foundation models on the LVIS and COCO benchmarks across different image quality settings. It is evident that foundation models are typically fine-tuned for a specific image resolution, which in turn yields the best quality and prompt alignment. In our experiments, DALL-E demonstrated a consistent improvement in performance with increasing image sizes, whereas Flux-Schnell and RealVisXL exhibited a less pronounced correlation between performance and image size, particularly between the MEDIUM and LARGE settings.
CLIPNeedleBothNeither0020202020404040406060606023.2323.2323.2323.2352.5252.5252.5252.5214.1514.1514.1514.1510.110.110.110.1Percentage (%)
Figure 11: Results from the human evaluation case study on the LVIS dataset. Among the queries with a clear preference, Needle was chosen in approximately 70% of cases, underscoring its practical effectiveness in retrieving relevant images.
Query Complexity Classifier: This section details the empirical investigation into the impact of our proposed Query Complexity Classifier on the overall efficiency and retrieval performance of Needle. The primary objective is to demonstrate how dynamically routing queries based on their predicted complexity can significantly reduce computational overhead without compromising retrieval effectiveness.
((a)) Comparison of Actual AP and Predicted AP using Needle Query Complexity Classifier
((b)) Comparison of Inference Speed and MAP using different baselines and benchmarks
Figure 12: Illustration of how Needle Query Complexity Classifier improves Needle inference speed
For these experiments, the query complexity classifier is implemented as a Linear Regression model, trained to predict the expected Average Precision (AP) of a given query when processed by preliminary retrieval methods. We specifically leverage features derived from two highly efficient preliminary methods: CLIP [53] and ALIGN [25]. The features utilized for this prediction, as detailed in Section C.4, include mean top-$K$ cosine similarity scores, top-$K$ inter-method overlap coefficients, and confidence deviation for both CLIP and ALIGN.
The model was trained on a combined dataset derived from the training splits of the Caltech256 [21] and LVIS [22] benchmarks. The ground-truth target for training was the actual AP achieved by the preliminary methods for each query. Figure 12(a) showcases the performance of our trained classifier in predicting AP for Caltech256.
The primary benefit of the query complexity classifier is its ability to dynamically route queries, thereby enhancing the inference speed of Needle. To quantify this improvement, we conducted experiments comparing the average inference speed and overall retrieval accuracy (Mean Average Precision, MAP) of Needle, both with and without the classifier, against several established baselines. The baselines include: CLIP, ALIGN, and BLIP+MiniLM. Our evaluation was performed across diverse benchmarks, including Caltech256, LVIS, and BDD100k [79]. Figure 12(b) illustrates a comprehensive comparison of average inference speed and retrieval accuracy for all evaluated baselines and configurations of Needle.
As one can observe from the results, employing our proposed query complexity classifier enables Needle to achieve significantly faster inference speeds with only a minimal or negligible loss in overall retrieval performance. This is primarily attributed to the classifier’s ability to accurately identify "simple" queries, for which the computationally intensive guide image generation process can be entirely bypassed. Furthermore, it is important to note that Needle’s efficiency is designed to progressively increase over time through its caching mechanism. As textual metadata for different images are extracted from user’s queries (C.3), the percentage of cache hits will naturally rise, further reducing the overhead of on-the-fly image generation and consequently enhancing inference speed.
E.5 Case Study
@@
Each human evaluator was asked to write five queries of their choice, with varying levels of complexity. For each query, we retrieved 20 relevant images from both CLIP and Needle, displaying the results side by side with randomized orientations for each query. Evaluators were then asked to select the more relevant results, choosing from four options: "Left is better," "Right is better," "Both are good enough," and "Neither is good enough". Then, we recorded the engine behind that option as the preferred engine.
@@
A total of 20 human evaluators participated in the study, contributing 100 queries. The queries ranged from simple ones like "Celery stew" to more complex descriptions such as "A person walking on the sidewalk of a river in Fall at Chicago.", Figure 11 presents the results of this user study. If we consider only the queries where one engine is preferred over the other, Needle is chosen as the better engine in almost 70% of cases.